OpenStudy (rjtrain1):

Tosha received a 9.5 for vault, a 9.1 for bars, a 9.625 for beam, and a 9.25 for her floor exercise at her last gymnastics meet. What was her combined (or total) score? 107.36 36.475 37.475 37.375 @Will.H

OpenStudy (will.h):

Total = 9.5 + 9.1 + 9.625 + 9.25 Total = ?? 1st add the constants together 9 + 9 + 9 + 9 = 36 now add the tenths together: 5 + 1 + 6 + 2 = 14 (we can't leave it like that we will add 1 to the constants and keep it to be 4 in the tenths) so the tenths now is 4 and the constant is 37 now add the hundredths: 2 + 2 = 4 now add the thousands: 5 + 5 = 10 the answer is 37.475